OpenStudy (anonymous):

Find the equation of a circle given that the points (8,0) and (0,6) are endpoints of a diameter

OpenStudy (anonymous):

Using the midpoint formula I know that part of the equation is: (x-4)^2 + (y-3)^2 =r^2 but I don't know r

OpenStudy (anonymous):

find the length of the diameter = sqrt((6-0)^2 + (0-8)^2) and radius is 1/2 of this

OpenStudy (anonymous):

D would equal 10, so the radius would be 5 not .5 though?

OpenStudy (anonymous):

correct - 5

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Oh, yes thought you were saying the radius was 1/2. Thank you

OpenStudy (anonymous):

yw